OpenStudy (anonymous):

what is the slope of the line passing through the points (2, 7) & (-1, 3)?

OpenStudy (shadowlegendx):

Remember the formula \[m = \frac{ y_{2} - y_{1} }{ x_{2} - x_{1} } \] m = slope

OpenStudy (shadowlegendx):

\[(x_{1},y_{1}) (x_{2},y_{2}) \]

OpenStudy (shadowlegendx):

(2, 7) (-1, 3)

OpenStudy (shadowlegendx):

Can you solve it now?

OpenStudy (shadowlegendx):

@mooniee ? Are you there?

OpenStudy (shadowlegendx):

\[m = \frac{ 3 - 7 }{ -1 - 2 }\]

OpenStudy (shadowlegendx):

The answer would be what? @mooniee ?

OpenStudy (shadowlegendx):

You should get: \[ -\frac{ 3 }{ 4 }\]